OpenStudy (anonymous):

f(x)=x^2-5x+4 and g(x)=3x+2 Evaluate f(g(4)) using the functions above.

OpenStudy (precal):

first do g(4)=3(4)+2 g(4)=12+2 = 14 f(g(4))=f(14)=(14)^2-5(4)+4=130

OpenStudy (anonymous):

yep that right

OpenStudy (kropot72):

The answer f(g(4)) = 130 is correct but in the substitution the term '-5(4)' should have been written -5(14)

OpenStudy (precal):

Yes, I stand corrected. f(14)=(14)^2-5(14)+4=130 thanks kropot72

OpenStudy (kropot72):

You're most welcome :)